OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -3a2 - b3 + 3c2 - 2b3 if a = 2, b = -1, and c = 3. Numerical Answers Expected!

TheSmartOne (thesmartone):

Can you plug in the values of a,b and c into the equation and see what you get. If you don't know then I'll help you.

TheSmartOne (thesmartone):

@shelbylynn

TheSmartOne (thesmartone):

Ok Lets first plug in a=2 So it will become -3(2)^2-b^3=3c^2-2b^3 which equals to -3(4)-b^3+3c^2-2b^3

TheSmartOne (thesmartone):

Now lets plug in b=-1 It will become -12-(-1)^3+3c^2-2(-1)^3 -1^3=-1x-1x-1=-1 So -12-(-1)+3c^2-2(-1) which equals to -12+1+3c^2+2 this equals to -9+3c^2

TheSmartOne (thesmartone):

Now plug in c=3 -9+3(3)^2 this equals to -9+3(9) which finally eqauls -9+27=18

TheSmartOne (thesmartone):

So your final answer is 18.

TheSmartOne (thesmartone):

@shelbylynn If you respond then I can assist you better.

OpenStudy (anonymous):

lol i was just cheking your math lol